OpenStudy (anonymous):

Help please with math! I need someone to help me find my answer step by step. A. solve a -9 = 20 B. Solve b - 9 > 20 C. How is solving the equation in part a similar to solving the inequality in part b? D. How are the solutions different?

OpenStudy (danjs):

ok

OpenStudy (danjs):

welcome to openstudy

OpenStudy (danjs):

A) a - 9 = 20

OpenStudy (danjs):

when you solve equations, you can do operations on them, but what you do to one side of the equation, you have to do to the other side.. try adding 9 to both sides of that equation a - 9 = 20

OpenStudy (anonymous):

a = 29

OpenStudy (danjs):

yep, you got that one

OpenStudy (danjs):

You can do the same thing with addition and subtraction with inequalities.... b - 9 > 20 ...........try adding 9 to this one on both sides also

OpenStudy (anonymous):

29

OpenStudy (danjs):

b > 29, remember to keep all the information

OpenStudy (danjs):

so for part C, what did you do to both part a and part b, that was similar

OpenStudy (anonymous):

I subtract both of them and i got 20. Like 29- 9 = 20.